# Model-free High Dimensional Mediator Selection with False Discovery Rate Control

### v1

@@ 5 Discussions
$\bm{\alpha}=(\,\underbrace{a,\ldots,a}_{9},\ \underbrace{0.7a,\ldots,0.7a}_{21},\ \underbrace{0,\ldots,0}_{p-30}\,)^{\top}$, $\bm{\beta}=(\,\underbrace{0,\ldots,0}_{9},\ \underbrace{0.7b,\ldots,0.7b}_{21},\ \underbrace{b,\ldots,b}_{10},\ \underbrace{0,\ldots,0}_{p-40}\,)^{\top}$.
In the simulation setting, we vary the correlation structure among mediators by setting the correlation coefficient $\rho\in\{0,0.1,\dots,0.9\}$. To examine the impact of signal strength in path a and path b, we vary $a\in\{0.1,0.2,0.3,0.4,0.5\}$ and $b\in\{0.1,0.2,0.3,0.4,0.5\}$. The number of potential mediators is set to vary across $p=100,200,\dots,500$ to evaluate performance in different dimensional settings. We fix $\gamma=1$,$\bm{\eta}_{1}=(1,\cdots,1)^{\top}$, $\eta_{2}=1$ and $\sigma=0.4$.
@@
In the simulation setting, we vary the correlation structure among mediators by setting the correlation coefficient $\rho\in\{0.1,0.3,0.5\}$. To examine the impact of signal strength in path a and path b, we vary $a\in\{0.4,0.6\}$ and $b\in\{0.5,1\}$. The number of potential mediators is set to be $p=100$. We fix $\gamma=1$,$\bm{\eta}_{1}=(0,\cdots,0)^{\top}$, and $\sigma=0.4$.
@@
$\bm{\alpha}=(\underbrace{a,\ldots,a}_{9},\ \underbrace{0.7a,\ldots,0.7a}_{21},\ \underbrace{0,\ldots,0}_{p-30}).$, $\bm{\beta}=(\underbrace{0,\ldots,0}_{9},\ \underbrace{b,\ldots,b}_{21},\ \underbrace{0.7b,\ldots,0.7b}_{9},\ \underbrace{0,\ldots,0}_{p-39}).$ In the simulation setting, we vary the correlation structure among mediators by setting the correlation coefficient $\rho\in\{0.1,0.3\}$. To examine the impact of signal strength in path a and path b, we vary $a\in\{0.2,0.5\}$ and $b\in\{0.5,1\}$. The number of potential mediators is set to be $p=400$. We fix $\gamma=1$,$\bm{\eta}_{1}=(0,\cdots,0)^{\top}$, and $\sigma=0.4$.
@@
Figure 2 summarizes the empirical results for non-linear models with interactions (Setting 1) and models where mediators undergo a cosine transformation. (Setting 2). In Setting 1, we observe that GKMS consistently achieves the highest power across all examined conditions, particularly under low to moderate correlation levels ($\rho\leq 0.5$). As the correlation among mediators increases, the power of all methods declines; however, GKMS demonstrates a more gradual decrease and remains superior throughout. When varying the signal strengths $a$ (path-a) and $b$ (path-b), GKMS continues to dominate in terms of power, while maintaining FDR well below or near the nominal level of 0.1. In contrast, eHIMA often fails to control FDR, particularly under weak signals or high correlation, and shows substantially reduced power.
@@
Figure 2: Simulation results for the empirical power and false discovery rate (FDR) with sample size $n=1000$ are shown for correlated mediators with a defined correlation structure across 100 replications. The results are stratified by the effect size of the correlation structure ($\rho$), the number of potential mediators (p), the effect size of path-a (a), and the effect size of path-b (b) in the non-linear models with Setting 1 in the upper, Setting in the lower.

### v2

@@ 5 Discussions
$\bm{\alpha}=(\,\underbrace{a,\ldots,a}_{9},\ \underbrace{0.7a,\ldots,0.7a}_{21},\ \underbrace{0,\ldots,0}_{p-30}\,)^{\top}$,
$\bm{\beta}=(\,\underbrace{0,\ldots,0}_{9},\ \underbrace{0.7b,\ldots,0.7b}_{21},\ \underbrace{b,\ldots,b}_{10},\ \underbrace{0,\ldots,0}_{p-40}\,)^{\top}$.
In the simulation setting, we vary the correlation structure among mediators by setting the correlation coefficient $\rho\in\{0,0.1,\dots,0.9\}$. To examine the impact of signal strength in path a and path b, we vary $a\in\{0.1,0.2,0.3,0.4,0.5\}$ and $b\in\{0.1,0.2,0.3,0.4,0.5\}$. The number of potential mediators is set to vary across $p=100,200,\dots,500$ to evaluate performance in different dimensional settings. We fix $\gamma=1$$\bm{\eta}_{1}=(1,\cdots,1)^{\top}$, $\eta_{2}=1$ and $\sigma=0.4$.
@@
In the simulation setting, we vary the correlation structure among mediators by setting the correlation coefficient $\rho\in\{0.1,0.3,0.5\}$. To examine the impact of signal strength in path a and path b, we vary $a\in\{0.4,0.6\}$ and $b\in\{0.5,1\}$. The number of potential mediators is set to be $p=100$. We fix $\gamma=1$$\bm{\eta}_{1}=(0,\cdots,0)^{\top}$, and $\sigma=0.4$.
@@
$\bm{\alpha}=(\underbrace{a,\ldots,a}_{9},\ \underbrace{0.7a,\ldots,0.7a}_{21},\ \underbrace{0,\ldots,0}_{p-30}).$,
$\bm{\beta}=(\underbrace{0,\ldots,0}_{9},\ \underbrace{b,\ldots,b}_{21},\ \underbrace{0.7b,\ldots,0.7b}_{9},\ \underbrace{0,\ldots,0}_{p-39}).$ In the simulation setting, we vary the correlation structure among mediators by setting the correlation coefficient $\rho\in\{0.1,0.3\}$. To examine the impact of signal strength in path a and path b, we vary $a\in\{0.2,0.5\}$ and $b\in\{0.5,1\}$. The number of potential mediators is set to be $p=400$. We fix $\gamma=1$$\bm{\eta}_{1}=(0,\cdots,0)^{\top}$, and $\sigma=0.4$.
@@
Figure C1: Simulation results for the empirical power and false discovery rate (FDR) with sample size $n=1000$ are shown for correlated mediators with a defined correlation structure across 100 replications. The results are stratified by the effect size of the correlation structure ($\rho$), the number of potential mediators (p), the effect size of path-a (a), and the effect size of path-b (b) in the non-linear models with Setting 1 in the upper, Setting in the lower.
Figure C1 summarizes the empirical results for non-linear models with interactions (Setting 1) and models where mediators undergo a cosine transformation. (Setting 2). In Setting 1, we observe that GKMS consistently achieves the highest power across all examined conditions, particularly under low to moderate correlation levels ($\rho\leq 0.5$). As the correlation among mediators increases, the power of all methods declines; however, GKMS demonstrates a more gradual decrease and remains superior throughout. When varying the signal strengths $a$ (path-a) and $b$ (path-b), GKMS continues to dominate in terms of power, while maintaining FDR well below or near the nominal level of 0.1. In contrast, eHIMA often fails to control FDR, particularly under weak signals or high correlation, and shows substantially reduced power.
@@
Appendix D.1: Structural MRI-Derived Measures
Appendix D.2: Cognitive Outcome Measures